OpenStudy (anonymous):

Solve each equation or inequality. If there is no solution, write no solution. 50. |n + 2| = 4 52. |x + 3| = -2 56. 4 |k + 5| > 8

hartnn (hartnn):

hint: if u have |a|=b the u can write a=b OR a=-b

OpenStudy (anonymous):

.........what

hartnn (hartnn):

for example, if u have |n + 2| = 4 you can write two equation and solve them individually, n+2 =4 OR n+2 =-4 can u solve these equations ?

OpenStudy (anonymous):

no I really don't understand this

hartnn (hartnn):

which part ? i think u can solve this to get n n+2=4 n=?

OpenStudy (anonymous):

2

hartnn (hartnn):

correct! now solve this to get other value of n n+2=-4

OpenStudy (anonymous):

-6

hartnn (hartnn):

so the two values , u got for |n + 2| = 4 are n=2 or -6 got this ?

OpenStudy (anonymous):

yeah but why is there two

hartnn (hartnn):

is your question., why there are 2 solutions or is it why is n=2 ?

OpenStudy (anonymous):

why are there two solutions?

hartnn (hartnn):

because, for example |a| =4 u can put a =4 or a=-4 so most of the absolute value equations do have 2 solutions. |....| makes the negative *minus* go away hence, |-4| also equals 4.

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhh okay

hartnn (hartnn):

AND |...| can never be a negative valued. so what can u say about question 52. ?

OpenStudy (anonymous):

-5?

hartnn (hartnn):

|x + 3| = -2 but |..*anything*..| can never be negative. so, |x+3| can never equal -2 hence, no solution, got this ?

hartnn (hartnn):

because '|....| makes the negative *minus* go away'

OpenStudy (anonymous):

ohh okay because since the | |'s make it not negative it wouldn't equal -2

hartnn (hartnn):

yes, |x+3| will always be positive for ANY value of x.

hartnn (hartnn):

try -5, |-5+3| = |-2| = ?

OpenStudy (anonymous):

wait what? i'm confused again

hartnn (hartnn):

as i said, |-a| = a so |-2| = 2 hence its not -2 , just trying to explain, why -5 is not the solution....

OpenStudy (anonymous):

ohhhhhhhh okay i didn't understand what you were asking. what about the last problem though? it's confusing because in the beginning there's a number and at the end instead of an equal sign there's a >

hartnn (hartnn):

4 |k + 5| > 8 first isolate the absolute sign, by dividing by 4 ,what u get ?

OpenStudy (anonymous):

what do i divide by 4? the 8?

hartnn (hartnn):

both sides of the equation

OpenStudy (anonymous):

so would the first 4 cancel out ? and then since id divide 4 and 8 it'd be 2?

hartnn (hartnn):

thats correct.

OpenStudy (anonymous):

so then it would be |k + 5| > 2??

hartnn (hartnn):

yes. now when |x|>a means x>a and x<-a so what about |k+5|>2 ?

OpenStudy (anonymous):

k - 5 >2 ?

hartnn (hartnn):

why k-5 ?

OpenStudy (anonymous):

i meant + sorry

hartnn (hartnn):

thats one eqution, solve it.

OpenStudy (anonymous):

would it just be 5k? i dont get it

hartnn (hartnn):

k+5 > 2 subtracting 5 from both sides k+5-5 > 2-5 k> -3 got this ?

OpenStudy (anonymous):

ohhh i didn't know i was supposed to subtract

hartnn (hartnn):

since k was ADDED to 5 , to isolate k, we need to SUBTRACT 5. now solve second equation k+5< -2

hartnn (hartnn):

again by subtracting 5 from both sides

OpenStudy (anonymous):

the answer would just be k> -3 then right? since that's what's left after you subtract the 5 from both sides?

hartnn (hartnn):

k < -2-5 k< -7

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

so after k> -2 you have to subtract 5 from the -2 again?

hartnn (hartnn):

k>-2 is one solution to the original equation, now u also get k< -7

hartnn (hartnn):

so overall -7 > k > -2

OpenStudy (anonymous):

ohhh okay thanks

hartnn (hartnn):

hope u understand a bit.... welcome :)

OpenStudy (anonymous):

50. n = -6 and n = 2 52. False cannot be solved 56. Complicated, not sure I'd be able to explain it to you